# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v3

@@ 3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination at step $T$. Then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (silver_general_2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network.
[…]
4 Experiments
4.1 Benchmark experiments
We benchmark AlphaDE and baselines on a suite of eight in-silicon protein engineering tasks (ren_proximal_2022). These tasks involve extensive protein applications such as biosensor design and industrial enzyme renovation. More descriptions about the eight benchmark tasks are provided in Appendix A. Following previous works (ren_proximal_2022; wang_self-play_2023; qiu_tree_2024), we simulate the ground-truth fitness landscape of protein with an oracle model TAPE (rao_evaluating_2019) to replace wet laboratory measurements.
We compare AlphaDE with various baselines including LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX (one version with a CNN model and another with a MuFacNet model), and EvoPlay for running. We run AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES with the implementations in FLEXS (sinai_adalead_2020), following the same benchmark setting, such as the starting sequence and fitness oracle as Ren et al. (ren_proximal_2022). We average the results of AlphaDE and all baselines over 5 independent trials. For all methods, we set the same oracle budget of 1000 queries. For AlphaDE, here we use the pretrained ESM2-series models (lin_evolutionary-scale_2023) for fine-tuning due to its strong performance, wide adoption, and availability in various model sizes, which is convenient for our study. Specifically, in the standard AlphaDE, we use ESM2-35M for fine-tuning due to its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence to evolve is the sequence with the lowest protein fitness. The benchmark results for eight tasks are summarized in Table 1, which includes mean values and standard deviations.
@@ 4.1 Benchmark experiments
We also test AlphaDE’s compatibility with other protein language models such as ProtBert (elnaggar_prottrans_2022) (model size 420M) and ESM-1b (rives_biological_2021) (model size 650M), which are trained with the standard BERT (devlin_bert_2019) architecture using the UniRef100 (suzek_uniref_2007) database and with the optimized RoBERTa (liu_roberta_2019) architecture using the UniRef50 (suzek_uniref_2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. Generally, AlphaDE with the two different protein language models achieves superior performance, demonstrating AlphaDE’s compatibility. On the other hand, AlphaDE with ProtBert and ESM-1b perform differently in some tasks, suggesting the evolution ability of protein language models varies in different protein families. Meanwhile, the fine-tuning step is the key to boosting performance, as directly using the pretrained versions of models leads to huge fitness loss.
@@
fine-tuned ProtBert (elnaggar_prottrans_2022) 3.09 7.66 0.49 7.66 0.00 -0.01 0.41 2.71
@@
fine-tuned ESM-1b (rives_biological_2021) 3.09 16.97 0.49 7.85 0.03 0.01 0.60 1.49
@@
4.2 Few-Shot Fine-Tuning Experiments
One of the biggest advantages of large language models is that they are few-shot learners (brown_language_2020). Here we investigate whether this advantage is held in the protein language model for directed evolution tasks. We use the ESM2-35M model as the base model for few-shot fine-tuning and fine-tune it with different numbers of protein sequences including 16, 32, 64, 128, 256, 512, and 1024. These protein sequences for few-shot fine-tuning are randomly sampled from the whole dataset and are used to fine-tune ESM2-35M with 3 epochs. The results of AlphaDE with different fine-tuned ESM2-35M models of different sequence numbers are given in Figure 2. We see that, even with dozens of protein sequences, the evolution ability of AlphaDE could be greatly improved, and the performance increases with the number of fine-tuning protein sequences.
@@ 4.2 Few-Shot Fine-Tuning Experiments
Although AlphaDE does not utilize fitness labels of protein sequences in the fine-tuning and MCTS inference steps, we additionally test whether the protein fitness influences AlphaDE’s performance by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3. The curves are very similar to Figure 2, showing that AlphaDE’s evolution capability is not affected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly wet measurement assays.
We also show that AlphaDE works with the sequences from homology searching such as HMMER (potter_hmmer_2018) in Appendix J, where homology sequences are retrieved from a database.
@@ 4.4 Condensing Protein Sequence Space
The origin and evolution of protein folds remain a fundamental challenge in biology (chothia_how_1997; levitt_nature_2009), with methods for exploring evolutionary trajectories still underdeveloped. Leveraging AlphaDE’s evolutionary capability, we address this through condensing the avGFP sequence space by evolving an incomplete, non-folding sequence into a functional, folded protein. Inspired by hypotheses that proteins evolve from small random peptides into complex folded structures (nepomnyachiy_complex_2017; kolodny_bridging_2021), we keep the chromophore and $\beta$-barrel residues of avGFP, mask half the remaining sequence, and use AlphaDE to optimize the predicted fluorescence fitness to recover protein function. If AlphaDE recovers the predicted fluorescence function with fewer amino acid residues than the wild-type avGFP sequence, it successfully condenses the protein sequence space through computational evolution. Fluorescence intensity is predicted using an ESM-1b-based oracle landscape simulator (ren_proximal_2022), where character ‘-’ denotes residue deletion. The computational evolution process is illustrated in Figure 5.
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (abramson_accurate_2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the comparison to AlphaDE without the fine-tuning step and the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ Appendix A Descriptions of Benchmark Tasks
Here we briefly describe the eight benchmark tasks constructed by Ren et al. (ren_proximal_2022).
@@
  • TEM-1 $\beta$-Lactamase (TEM). TEM-1 $\beta$-Lactamase protein resisting penicillin antibiotics in E.coli is widely studied to understand the mutational effect and fitness landscape (bershtein_robustnessepistasis_2006; jacquier_capturing_2013). Here we optimize the thermodynamic stability in the protein sequence space with a size of $20^{286}$.
@@ C.3 Value Network Settings in AlphaDE
The value network of AlphaDE is with the same network architecture as in EvoPlay (wang_self-play_2023), which consists of 4 convolutional neural network layers and 2 dense layers. As the same network architecture works well in AlphaDE, we do not modify the value network architecture. The only difference is that we do not specify the activation function in the output layer, as we found that the predicted fitness value of different tasks has different value ranges, and we do not restrict the output range of the value network.
@@ Appendix E Fitness Oracles
The use of TAPE and ESM-1b oracles follows the standard evaluation procedure in previous works (ren_proximal_2022; wang_self-play_2023; qiu_tree_2024). The ESM-1b oracle is used in the avGFP condensing experiment as it supports some special characters, including the token ‘-’. We regard the character token ‘-’ as an amino acid residue deletion and utilize ESM-1b to predict the fitness of the condensed avGFP sequences. These oracle models and their download script can be obtained from GitHub.
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
We conduct the benchmark experiments with ESM-1b oracle taken from ren_proximal_2022. Results are shown in Table 6. AlphaDE again beats various competitive baselines in the benchmark setting of EMS-1b oracle.
@@ Appendix I Diversity of Evolved Sequences by AlphaDE
In the benchmark experiments, almost all the fine-tuned sequence datasets are from deep mutational scanning (DMS), which are not always available for most proteins. Therefore, in this section, we introduce the homology searching technique to construct the dataset of homologous sequences for the fine-tuning step in AlphaDE. Next, we use avGFP as an example. Specifically, we use the phmmer homology searching procedure of the biosequence analysis tool HMMER (potter_hmmer_2018) to find homologous sequences of the starting weakest sequence in the avGFP task. We use the default settings of phmmer to search the SwissProt database, UniProt database, Reference Proteomes database, and PDB database. After filtering sequences with the same length as avGFP and removing three duplicate sequences in the DMS dataset, we found 236 unique sequences to construct the avGFP phmmer dataset, which has no overlap with the avGFP DMS dataset. Then we follow the standard AlphaDE fine-tuning step on the avGFP phmmer dataset and perform the MCTS step with the fine-tuned model (here $c$ is set at 1.0 for a better performance). Results are shown in Table 12 below. We see that, AlphaDE, which uses the phmmer homologous sequence dataset for fine-tuning, also achieves superior performance.
@@
When initializing the deleted avGFP sequences, we keep the $\beta$-barrel residues and the chromophore-related residues. The $\beta$-barrel residues involve the residue 1 to residue 38. The key chromophore residues are residue 65, 66, and 67 (hayes_simulating_2025), and we set the chromophore-related residues to be residue 58 to residue 74. Then we delete half of the remaining residues, which results in a starting deleted sequence with length 146. In contrast, the wild-type sequence has a length of 238 residues. The wild-type avGFP sequence, the starting deleted sequence, and the final filtered ccGFP1-5 sequences are given in Table 13. When filtering, the amino acid sequences of ccGFP1-5 are fed into AlphaFold 3 server to predict their structures. Then we align these structures with the wild-type structure (PDB ID: 1EMA) by PyMol to calculate the RMSD distances. Although this is a computational proof-of-concept task, it shows the great potential of AlphaDE for different purposes with directed evolution.

### v4

@@ 3.3 MCTS with Fine-Tuned Protein Language Model
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination at step $T$. Then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (Silver et al., 2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network.
[…]
4 Experiments
4.1 Benchmark experiments
@@
fine-tuned ProtBert (Elnaggar et al., 2022) 3.09 7.66 0.49 7.66 0.00 -0.01 0.41 2.71
@@
fine-tuned ESM-1b (Rives et al., 2021) 3.09 16.97 0.49 7.85 0.03 0.01 0.60 1.49
@@
We benchmark AlphaDE and baselines on a suite of eight in-silicon protein engineering tasks (Ren et al., 2022). These tasks involve extensive protein applications such as biosensor design and industrial enzyme renovation. More descriptions about the eight benchmark tasks are provided in Appendix A. Following previous works (Ren et al., 2022; Wang et al., 2023; Qiu et al., 2024), we simulate the ground-truth fitness landscape of protein with an oracle model TAPE (Rao et al., 2019) to replace wet laboratory measurements.
We compare AlphaDE with various baselines including LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX (one version with a CNN model and another with a MuFacNet model), and EvoPlay for running. We run AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES with the implementations in FLEXS (Sinai et al., 2020), following the same benchmark setting, such as the starting sequence and fitness oracle as Ren et al. (Ren et al., 2022). We average the results of AlphaDE and all baselines over 5 independent trials. For all methods, we set the same oracle budget of 1000 queries. For AlphaDE, here we use the pretrained ESM2-series models (Lin et al., 2023) for fine-tuning due to its strong performance, wide adoption, and availability in various model sizes, which is convenient for our study. Specifically, in the standard AlphaDE, we use ESM2-35M for fine-tuning due to its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence to evolve is the sequence with the lowest protein fitness. The benchmark results for eight tasks are summarized in Table 1, which includes mean values and standard deviations.
We also test AlphaDE’s compatibility with other protein language models such as ProtBert (Elnaggar et al., 2022) (model size 420M) and ESM-1b (Rives et al., 2021) (model size 650M), which are trained with the standard BERT (Devlin et al., 2019) architecture using the UniRef100 (Suzek et al., 2007) database and with the optimized RoBERTa (Liu et al., 2019) architecture using the UniRef50 (Suzek et al., 2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. Generally, AlphaDE with the two different protein language models achieves superior performance, demonstrating AlphaDE’s compatibility. On the other hand, AlphaDE with ProtBert and ESM-1b perform differently in some tasks, suggesting the evolution ability of protein language models varies in different protein families. Meanwhile, the fine-tuning step is the key to boosting performance, as directly using the pretrained versions of models leads to huge fitness loss.
4.2 Few-Shot Fine-Tuning Experiments
One of the biggest advantages of large language models is that they are few-shot learners (Brown et al., 2020). Here we investigate whether this advantage is held in the protein language model for directed evolution tasks. We use the ESM2-35M model as the base model for few-shot fine-tuning and fine-tune it with different numbers of protein sequences including 16, 32, 64, 128, 256, 512, and 1024. These protein sequences for few-shot fine-tuning are randomly sampled from the whole dataset and are used to fine-tune ESM2-35M with 3 epochs. The results of AlphaDE with different fine-tuned ESM2-35M models of different sequence numbers are given in Figure 2. We see that, even with dozens of protein sequences, the evolution ability of AlphaDE could be greatly improved, and the performance increases with the number of fine-tuning protein sequences.
Although AlphaDE does not utilize fitness labels of protein sequences in the fine-tuning and MCTS inference steps, we additionally test whether the protein fitness influences AlphaDE’s performance by fine-tuning on the sampled sequences from the bottom 20% of data with the lowest fitness. The results of fine-tuning with low-fitness data are plotted in Figure 3. The curves are very similar to Figure 2, showing that AlphaDE’s evolution capability is not affected by fitness values during fine-tuning. This suggests applicability in real-world settings where only sequences are available, without costly wet measurement assays. We also show that AlphaDE works with the sequences from homology searching such as HMMER (Potter et al., 2018) in Appendix J, where homology sequences are retrieved from a database.
@@ 4.4 Condensing Protein Sequence Space
The origin and evolution of protein folds remain a fundamental challenge in biology (Chothia and Gerstein, 1997; Levitt, 2009), with methods for exploring evolutionary trajectories still underdeveloped. Leveraging AlphaDE’s evolutionary capability, we address this through condensing the avGFP sequence space by evolving an incomplete, non-folding sequence into a functional, folded protein. Inspired by hypotheses that proteins evolve from small random peptides into complex folded structures (Nepomnyachiy et al., 2017; Kolodny et al., 2021), we keep the chromophore and $\beta$-barrel residues of avGFP, mask half the remaining sequence, and use AlphaDE to optimize the predicted fluorescence fitness to recover protein function. If AlphaDE recovers the predicted fluorescence function with fewer amino acid residues than the wild-type avGFP sequence, it successfully condenses the protein sequence space through computational evolution. Fluorescence intensity is predicted using an ESM-1b-based oracle landscape simulator (Ren et al., 2022), where character ‘-’ denotes residue deletion. The computational evolution process is illustrated in Figure 5.
We conduct 10 condensing trials and select the top 2 sequences from each trial with the highest fitness while keeping the number of deleted residues no less than 50. Then we use AlphaFold 3 (Abramson et al., 2024) to predict structures of these 20 sequences and align their structures with the wild-type avGFP (PDB ID: 1EMA). Finally, we filter the top 5 sequences (ccGFP1-5) with the smallest RMSD to the wild-type structure, as in Figure 5. As we see, AlphaDE successfully evolves the truncated, non-folding sequence to predictedly functional, folded proteins with fewer residues than the wild-type avGFP. More details, including the comparison to AlphaDE without the fine-tuning step and the amino acid residues of ccGFP1-5’s protein sequences, can be referred to in Appendix K.
@@ Appendix A Descriptions of Benchmark Tasks
Here we briefly describe the eight benchmark tasks constructed by Ren et al. (Ren et al., 2022).
@@
  • TEM-1 $\beta$-Lactamase (TEM). TEM-1 $\beta$-Lactamase protein resisting penicillin antibiotics in E.coli is widely studied to understand the mutational effect and fitness landscape (Bershtein et al., 2006; Jacquier et al., 2013). Here we optimize the thermodynamic stability in the protein sequence space with a size of $20^{286}$.
@@ C.3 Value Network Settings in AlphaDE
The value network of AlphaDE is with the same network architecture as in EvoPlay (Wang et al., 2023), which consists of 4 convolutional neural network layers and 2 dense layers. As the same network architecture works well in AlphaDE, we do not modify the value network architecture. The only difference is that we do not specify the activation function in the output layer, as we found that the predicted fitness value of different tasks has different value ranges, and we do not restrict the output range of the value network.
@@ Appendix E Fitness Oracles
The use of TAPE and ESM-1b oracles follows the standard evaluation procedure in previous works (Ren et al., 2022; Wang et al., 2023; Qiu et al., 2024). The ESM-1b oracle is used in the avGFP condensing experiment as it supports some special characters, including the token ‘-’. We regard the character token ‘-’ as an amino acid residue deletion and utilize ESM-1b to predict the fitness of the condensed avGFP sequences. These oracle models and their download script can be obtained from GitHub.
@@ Appendix G Benchmark Experiments with ESM-1b Oracle
We conduct the benchmark experiments with ESM-1b oracle taken from Ren et al. (2022). Results are shown in Table 6. AlphaDE again beats various competitive baselines in the benchmark setting of EMS-1b oracle.
@@ Appendix I Diversity of Evolved Sequences by AlphaDE
In the benchmark experiments, almost all the fine-tuned sequence datasets are from deep mutational scanning (DMS), which are not always available for most proteins. Therefore, in this section, we introduce the homology searching technique to construct the dataset of homologous sequences for the fine-tuning step in AlphaDE. Next, we use avGFP as an example. Specifically, we use the phmmer homology searching procedure of the biosequence analysis tool HMMER (Potter et al., 2018) to find homologous sequences of the starting weakest sequence in the avGFP task. We use the default settings of phmmer to search the SwissProt database, UniProt database, Reference Proteomes database, and PDB database. After filtering sequences with the same length as avGFP and removing three duplicate sequences in the DMS dataset, we found 236 unique sequences to construct the avGFP phmmer dataset, which has no overlap with the avGFP DMS dataset. Then we follow the standard AlphaDE fine-tuning step on the avGFP phmmer dataset and perform the MCTS step with the fine-tuned model (here $c$ is set at 1.0 for a better performance). Results are shown in Table 12 below. We see that, AlphaDE, which uses the phmmer homologous sequence dataset for fine-tuning, also achieves superior performance.
@@
When initializing the deleted avGFP sequences, we keep the $\beta$-barrel residues and the chromophore-related residues. The $\beta$-barrel residues involve the residue 1 to residue 38. The key chromophore residues are residue 65, 66, and 67 (Hayes et al., 2025), and we set the chromophore-related residues to be residue 58 to residue 74. Then we delete half of the remaining residues, which results in a starting deleted sequence with length 146. In contrast, the wild-type sequence has a length of 238 residues. The wild-type avGFP sequence, the starting deleted sequence, and the final filtered ccGFP1-5 sequences are given in Table 13. When filtering, the amino acid sequences of ccGFP1-5 are fed into AlphaFold 3 server to predict their structures. Then we align these structures with the wild-type structure (PDB ID: 1EMA) by PyMol to calculate the RMSD distances. Although this is a computational proof-of-concept task, it shows the great potential of AlphaDE for different purposes with directed evolution.